LLOP corporation just paid 4$ dividend per share, you expect the dividend to grow 8% for the next 2 years and expect to sell the stock at $50 at the end of year 2. What is the maximum prie you would pay to buy the stock? the required rate of return is 15%.
Answer #1

Year 1 dividend = 4 * 1.08 = 4.32

Year 1 dividend = 4.32 * 1.08 = 4.6656

Maximum price = 4.32 / (1 + 0.15)1 + 4.6656 / (1 + 0.15)2 + 50 / (1 + 0.15)2

Maximum price = $45.09